A force 0f 100 N is directed north on a -20 µC point charge. What is the magnitude and direction of the electric field at this point?


1
Expert's answer
2022-05-22T14:37:55-0400
"F=qE"

"E=\frac{F}{q}=\frac{100\:\rm N}{20*10^{-6}\:\rm N/C}=5*10^6\:\rm N"

The direction of the electric field at this point is toward south.